OpenStudy (anonymous):

Find the value of x for which f(x) = g(x). f(x) = sqrt(x) − 11, g(x) = 9 − x

OpenStudy (saifoo.khan):

\[\sqrt{x}-11 = 9-x\]

OpenStudy (anonymous):

you have to find the value of x

OpenStudy (saifoo.khan):

Righto.

OpenStudy (saifoo.khan):

Solve it.

OpenStudy (anonymous):

i tried and came up with something ridiculous

OpenStudy (anonymous):

remind me of the quadratic formula please

OpenStudy (saifoo.khan):

LOOL.

OpenStudy (saifoo.khan):

\[\sqrt{x}=20-x\]taking square on both sides:

OpenStudy (saifoo.khan):

\[x=(20-x)^2\]use foil to solve.

OpenStudy (anonymous):

still having trouble finding the answer. can you write out steps? im genuinely interested in figuring out this process. i have an exam next week